OpenStudy (anonymous):

How to convert a parabola in standard form into vertex form? One of my problems: f(x) = -3x+24x-51 Thanks!

OpenStudy (anonymous):

I have tried that site, but it doesn't give a direct step-by-step on how to solve the problem.

OpenStudy (anonymous):

-3x^2?

OpenStudy (anonymous):

Yes sorry!

OpenStudy (anonymous):

Thank you very much! This helped immensely.

OpenStudy (jdoe0001):

\(\bf -3(x^2-8x+16-16)-51\\ \quad \\\implies (-3\cdot x^2-(-3\cdot 8x)+(-3\cdot 16)-(-3\cdot 16))-51\)

OpenStudy (jdoe0001):

to complete the square, to get a perfect square trinomial, you're really just borrowing from zero, 0, so if you add whatever amount you need for the trinomial, you'd subtract that much